Substitition Method, Please help me the solutions
3x+4y=4
2x y=6
-y-(6+2x)
3x+4(6-2x)=4
3x+24-8x=4
5x=20/5=(4,2)


Answer by Alan3354(69443)

Substitition Method, Please help me the solutions
3x+4y=4
2x y=6
-y-(6+2x)
3x+4(6-2x)=4
3x+24-8x=4
5x=20/5=(4,2)
================
Is it all one problem? Separate the pairs.
----
2x y=6 No operator between terms.
-y-(6+2x) No equal sign.
5x=20/5=(4,2) Two equal signs.
-----------
Make an effort.